# Semantic Information Management in Low-Temperature Plasma Science and Technology with VIVO

### v1

@@ 3.2 Semantic Information Retrieval Using SPARQL
    “Which Dataset is obtained with DiagnosticMethod applied to diagnose PlasmaSource?”
@@
PREFIX plasma-o: <http://plasma-mds.org/ontology/>
@@
SELECT ?Person ?PersonLabel ?PlasmaSource ?Institute
@@
# Define the type (the class) of Person
@@
?Person plasma-o:isAffiliatedWith ?Institute .
?Person plasma-o:isExperiencedInDevice ?PlasmaSource .
@@
Listing 2 represents the SPARQL query for the second competency question. In this case, the query selects DiagnosticMethod and PlasmaSource instances with the labels “Cavity Ring-Down Spectroscopy” and “kINPen-sci”, respectively, to infer the related datasets (of class Dataset). The Dataset class is connected with DiagnosticMethod and PlasmaSource classes mainly through PlasmaStudy. Therefore, the query first infers the plasma studies that are related to the selected DiagnosticMethod and PlasmaSource using the relevant properties, then finally, it infers the datasets related to the inferred plasma studies. Similarly, the answer of this query is shown in Table 2, which returns two datasets.
Listing 2: SPARQL query for “Which Dataset is obtained with DiagnosticMethod applied to diagnose PlasmaSource?”
@@
PREFIX plasma-o: <http://plasma-mds.org/ontology/>
@@
SELECT ?Dataset ?DatasetLabel ?PlasmaStudy ?DiagnosticMethod ?PlasmaSource
@@
# Define Dataset and PlasmaStudy
@@
?PlasmaStudy rdf:type plasma-o:PlasmaStudy .
@@
?DiagnosticMethod rdfs:label "Cavity Ring-Down Spectroscopy (CRDS)"@en-US .
@@
?PlasmaStudy plasma-o:experimentUsesMethod ?DiagnosticMethod .
?PlasmaStudy plasma-o:hasPlasmaSource ?PlasmaSource .
?Dataset plasma-o:isOutputOf ?PlasmaStudy .
# Note: Dataset is connected to PlasmaStudy by the isOutputOf property,
@@
Table 2: SPARQL results for “Which Dataset is obtained with DiagnosticMethod applied to diagnose PlasmaSource?” (Listing 2).
Dataset DatasetLabel PlasmaStudy DiagnosticMethod PlasmaSource
@@ 4 Conclusion and Outlook
The present work introduces the initial foundation of a domain-specific ontology for low-temperature plasma science and technology, named Plasma-O, and its application using the VIVO software to build a knowledge graph for the LTP field, named Plasma-KG. By aligning Plasma-O with the existing VIVO ontology, we expand VIVO’s capabilities to include specialized applications like the semantic cataloging of plasma sources and datasets, and the ability to address complex research questions through SPARQL queries. These enhanced capabilities directly contribute to the adoption of the FAIR data principles in the LTP field and connects the often disjunct topics of current research information management, research data management and semantic knowledge management in an interoperable way. The improved discoverability and findability of complex data facilitated by Plasma-O and VIVO significantly enhance the Findability and Accessibility aspects of FAIR data principles. Moreover, the standardized, machine-readable format of ontologies and knowledge graphs inherently promotes Interoperability and Reusability of LTP knowledge and research data. The presented methods provide a generalizable platform for domain-specific semantics, which can also be used by third-party services, e.g. to annotate research assets such as patents and scientific literature [Aras et al. (2024)]. Further, this work aims to initiate related discussions within the LTP community, hopefully steering the future development of the ontology and knowledge graph.
Currently, Plasma-O models general concepts in the LTP field, which limits the types and complexity of competency questions that can be formulated. In future work, we intend to expand the ontology to cover fine-grained details of plasma studies, such as properties of devices and samples. This could be relevant not only for knowledge discovery but also during the experimentation stage of research. Consequently, we plan to integrate the ontology and knowledge graph into an RDM tool like Adamant [^9] https://github.com/plasma-mds/adamant [Chaerony Siffa et al. (2022)]. This would allow researchers to interact with the ontology, thus the knowledge graph, as early as the experimentation stage (e.g. while documenting the experiments in the electronic laboratory notebook), enabling them to receive suggestions or reuse existing metadata for their experiments.
@@
The VIVO software version 1.13 is used in this work and publicly available at the corresponding GitHub release page [^10] https://github.com/vivo-project/VIVO/releases/tag/vivo-1.13.0. The Plasma Ontology and the example SPARQL queries presented in this work are publicly available at the corresponding GitHub page  [^11] https://github.com/plasma-mds/plasma-ontology. The plasma ontology (version 0.5.0) and knowledge graph used in this work are available at: https://doi.org/10.5281/zenodo.13325226. The VIVO software platform containing the plasma knowledge graph is currently under preparation and will be openly available following an embargo at: https://vivo.plasma-mds.org.

### v3

@@ 3.2 Semantic Information Retrieval Using SPARQL
    “Which Dataset is obtained with Method applied to diagnose Plasma Source?”
@@
PREFIX plsmo: <https://plasma-mds.org/ontology/plasma-o/>
@@
PREFIX vivo: <http://vivoweb.org/ontology/core#>
@@
SELECT DISTINCT ?Person ?PersonLabel ?PlasmaSource ?Institute
@@
# Define the types of the involved entities
@@
?Institute rdf:type vivo:Institute .
?PlasmaSource rdf:type plsmo:PLSMO_C0005 .
@@
# Get Person and Institute instances that are connected with
# a predicate "affiliatedWith" which has IRI plsmo:PLSMO_R0003
?Person plsmo:PLSMO_R0003 ?Institute .
# Get Person and PlasmaSource instances that are connected with
# a predicate "experiencedIn" which has IRI plsmo:PLSMO_R0032
?Person plsmo:PLSMO_R0032 ?PlasmaSource .
@@
Similarly, Listing 2 represents the SPARQL query for the second competency question. In this case, the query selects Method and Plasma Source instances with the labels “Cavity Ring-Down Spectroscopy” and “kINPen-sci”, respectively, to infer the related datasets (of class Dataset). The Dataset class is connected with Method and Plasma Source classes mainly through Plasma Study. Therefore, the query first infers the plasma studies that are related to the selected Method and Plasma Source using the relevant properties, then finally, it infers the datasets related to the inferred plasma studies. Similarly, the answer of this query is shown in Table 3, which returns two datasets.
Listing 2: SPARQL query for “Which Dataset is obtained with Method applied to diagnose Plasma Source?”
@@
PREFIX plsmo: <https://plasma-mds.org/ontology/plasma-o/>
@@
SELECT DISTINCT ?Dataset ?DatasetLabel ?PlasmaStudy ?Method ?PlasmaSource
@@
# Define the types of the involved entities
@@
?PlasmaStudy rdf:type plsmo:PLSMO_C0024 .
?Method rdf:type plsmo:PLSMO_C0012 .
?PlasmaSource rdf:type plsmo:PLSMO_C0005 .
@@
?Method rdfs:label "Cavity Ring-Down Spectroscopy (CRDS)"@en-US .
@@
# Get instances of PlasmaStudy and Method connected
# via a predicate "usesMethod" which has IRI plsmo:PLSMO_R0105
?PlasmaStudy plsmo:PLSMO_R0105 ?Method .
# Get instances of PlasmaStudy and PlasmaSource connected
# via a predicate "hasPlasmaSource" which has IRI plsmo:PLSMO_R0039
?PlasmaStudy plsmo:PLSMO_R0039 ?PlasmaSource .
# Finally, get instances of Dataset and PlasmaStudy connected
# via a predicate "outputOf" which has IRI plsmo:PLSMO_R0062
?Dataset plsmo:PLSMO_R0062 ?PlasmaStudy .
# Note: Dataset is connected to PlasmaStudy by the outputOf property,
@@
Table 3: SPARQL results for “Which Dataset is obtained with Method applied to diagnose Plasma Source?” (Listing 2).
Dataset DatasetLabel PlasmaStudy Method PlasmaSource
@@ 4 Conclusion and Outlook
The present work introduces the initial foundation of a domain-specific ontology for low-temperature plasma science and technology, named Plasma-O, and its application using the VIVO platform to build a knowledge graph for the LTP field, named Plasma-KG. By aligning Plasma-O with the existing VIVO ontology, we expand VIVO’s capabilities to include specialized applications like the semantic cataloging of plasma sources and datasets, and the ability to address complex research questions through SPARQL queries. These enhanced capabilities directly contribute to the adoption of the FAIR data principles in the LTP field and connects the often disjunct topics of current research information management, research data management and semantic knowledge management in an interoperable way. The improved discoverability and findability of complex data facilitated by Plasma-O and VIVO significantly enhance the Findability and Accessibility aspects of FAIR data principles. Moreover, the standardized, machine-readable format of ontologies and knowledge graphs inherently promotes Interoperability and Reusability of LTP knowledge and research data. The presented methods provide a generalizable platform for domain-specific semantics, which can also be used by third-party services, e.g. to annotate research assets such as patents and scientific literature [Aras et al. (2024)]. Furthermore, the methods are readily adaptable to other adjacent applied research fields, in which domain-specific ontologies may have been developed. The potential for alignment with Plasma-O exists where relevant. For example, Plasma-O provides the basis to semantically integrate more generic metadata schemas like Plasma-MDS with more specific schemas, e.g. OpenPMD for particle simulations [Huebl et al. (2017)]. Moreover, Plasma-O allows to connect the concepts from LTP research with adjacent domains, e.g. material science via PMDco [Bayerlein et al. (2024)], and particle physics [Könnecke et al. (2015)]. Most importantly, this work aims to stimulate related discussions within the LTP community, hopefully steering the future development of the ontology and knowledge graph.
Currently, Plasma-O models general concepts in the LTP field, which limits the types and complexity of competency questions that can be formulated. In future work, we intend to expand the ontology to cover fine-grained details of plasma studies and experiments (see Fig. 1), such as Property of Target and Plasma, Reactive Species (e.g. $\mathrm{N_{2}^{+}}$ and $\mathrm{O_{2}^{-}}$) produced from the interaction between Plasma and Target (e.g. water), physical quantities such as ElectrodeLength and TreatmentDistance, etc. This could be relevant not only for knowledge discovery but also during the experimentation stage of research. Consequently, we plan to integrate the ontology and knowledge graph into an RDM tool like Adamant [^11] https://github.com/plasma-mds/adamant [Chaerony Siffa et al. (2022)]. This would allow researchers to interact with the ontology, thus the knowledge graph, as early as the experimentation stage (e.g. while documenting the experiments in the electronic laboratory notebook), enabling them to receive suggestions or reuse existing metadata for their experiments.
@@
The VIVO software platform version 1.15 is used in this work and publicly available at the corresponding GitHub release page [^12] https://github.com/vivo-project/VIVO/releases/tag/vivo-1.15.0. The Plasma Ontology and the example SPARQL queries presented in this work are publicly available at the corresponding GitHub page [^13] https://github.com/plasma-mds/plasma-ontology. The plasma ontology (version 0.7.0) and knowledge graph used in this work are available at: https://doi.org/10.5281/zenodo.14610064. The VIVO platform containing the plasma knowledge graph is openly available at: https://vivo.plasma-mds.org.